integrate the following integral
∫ (1/(7x-5)^3)  + (1/sqrt(5x-4)) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int [\frac{1}{(7x-5)^3}+\frac{1}{\sqrt{5x-4}}]dx
I = \int \frac{1}{(7x-5)^3}dx+\int \frac{1}{\sqrt{5x-4}}dx
I = \frac{1}{-2.7(7x-5)^2}dx+\int \frac{1}{\sqrt{5x-4}}dx
I = \frac{1}{-14(7x-5)^2}dx+ \frac{2\sqrt{5x-4}}{5}+c
I = \frac{-1}{14(7x-5)^2}dx+ \frac{2\sqrt{5x-4}}{5}+c